OpenStudy (anonymous):

integration using u sub of cos^2(sinx)cosx

hartnn (hartnn):

did you try u = sin x ?

OpenStudy (anonymous):

yup

hartnn (hartnn):

where are you stuck ? du = .. ?

OpenStudy (anonymous):

du=cosxdx, dx=du/cosx

hartnn (hartnn):

correct, go on

OpenStudy (anonymous):

used the trig identity that cos^2(u)= 1+cos(2u)/2

OpenStudy (anonymous):

canceled one of the cosx with the added du/cosx

OpenStudy (anonymous):

and got 1/2(1+cos(2u)

hartnn (hartnn):

and you still need to integrate that, isn't it ? with respect to 'u'

hartnn (hartnn):

and bdw, thats correct till there

hartnn (hartnn):

i am sure you know the integration of 1 are u stuck at finding integration of cos 2u ?

OpenStudy (anonymous):

no derived with repect to x cuz i came up with 1/2(x + sin(2u)/2+c, but should have been 1/2(du+sin(2sin(x)/2)

hartnn (hartnn):

1/2(u+ sin(2u)/2+c

hartnn (hartnn):

u/2 + (sin 2u)/4 + c now plug back in u = sin x

hartnn (hartnn):

\(\Large \int 1 du =u+c\)

OpenStudy (anonymous):

yup got it from there, thanks for the help bud

hartnn (hartnn):

welcome :)